OpenStudy (anonymous):

4/5 y + 2 =6 5 7 9 11

OpenStudy (dwelsher):

5

OpenStudy (anonymous):

solve by canceling out the fraction first

OpenStudy (dwelsher):

1.4/5y+2=6 2.4/5y=4 3.multiply both side by 5---4y=20 4sove for y

OpenStudy (anonymous):

EXAMPLE 2 3 [ --- x + 2 = 2 ] MULTIPLY WHOLE EQUATION BY 3 3 2X + 2 = 6 3 IS CANCELLED BECAUSE IT WAS A FRACTION 2X + 2 + (-2) = 6 + (-2 ) ADD ADDITIVE INVERSE 2X = 4 1/2 [ 2 X = 4 ] MULTIPLY BY RECIPROCAL X = 2